# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v3

@@ Abstract
In this work we systematically review the recent advancements in code processing with language models, covering 50+ models, 30+ evaluation tasks, 170+ datasets, and 700 related works. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also discuss code-specific features such as AST, CFG, and unit tests, along with their application in training code language models, and identify key challenges and potential future directions in this domain. We keep the survey open and updated on GitHub at https://github.com/codefuse-ai/Awesome-Code-LLM.
@@ 1 Introduction
Language modeling has advanced remarkably in recent years with the advent of pretrained Transformers (Vaswani et al., 2017) such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2018). As large language models (LLMs) scaled to hundreds of billions of parameters and started to display early signs of artificial general intelligence (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023), their applications have also transcended text processing. Pioneered by Codex (Chen et al., 2021b), LLMs have achieved impressive results in code processing, giving rise to commercial products such as GitHub Copilot111https://github.com/features/copilot and open-source multi-billion code models such as StarCoder (Li et al., 2023g) and Code LLaMA (Rozière et al., 2023).
The application of pretrained Transformers in code processing, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for code in this work, covering 50+ models, 30+ downstream tasks, 170+ datasets, and 700 related works. We break down different categories of code language models, ranging from colossal models trained on general domains to tiny models trained specifically for code understanding or generation. We emphasize on the relations and differences between such models, and highlight the integration of code-specific features, such as abstract syntax trees or data flows, into language models, as well as the latest techniques adapted from NLP.
Related to our work, we are aware of several surveys on similar topics, with three works concurrent to us (Hou et al., 2023; Zheng et al., 2023b; She et al., 2023). These works, however, focus either on NLP side (Zan et al., 2023; Xu & Zhu, 2022) or SE side (Niu et al., 2023; Hou et al., 2023; Zheng et al., 2023b; She et al., 2023), and do not cover models, tasks, and challenges from the other side. For example, Zan et al. (2023) focus on LLMs for text-to-code generation, while giving little discussion of other evaluation tasks in software engineering community. Hou et al. (2023) and She et al. (2023), in contrast, comprehensively review works from SE venues such as ASE and ICSE, but cite only a handful of works from deep learning and NLP venues such as ACL, EMNLP, NeurIPS, and ICLR.
@@
where level=2text width=4.5em,font=,
@@
[Code LMs
[Raw LM
[LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2022), GPT-NeoX (Black et al., 2022),
@@
LLaMA 2 (Touvron et al., 2023b), Phi-1.5 (Li et al., 2023h),
text width=27em,font=]
[Adapted LM
[Codex (Chen et al., 2021b), PaLM Coder (Chowdhery et al., 2022), Minerva (Lewkowycz et al., 2022),
PaLM 2* (Anil et al., 2023), Code LLaMA (Rozière et al., 2023),
text width=27em,font=]
[Specialized LM
[Encoder
@@
text width=20.5em]
[Decoder
[CLM
@@
text width=16.8em]
[FIM
[InCoder (Fried et al., 2023), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023g),
text width=16.8em]
[UniLM
@@
text width=20.5em]
[Encoder-Decoder
@@
CodeT5+ (Wang et al., 2023d),
text width=20.5em]
@@
Finetuning
[Instruction Finetuning
[WizardCoder (Luo et al., 2023), PanGu-Coder2 (Shen et al., 2023), OctoCoder (Muennighoff et al., 2023),
MFTCoder (Liu et al., 2023b),
text width=20.5em]
[Reinforcement Learning
@@
text width=20.5em]
@@
Thus, building on these works, we endeavor to unite the perspectives from both communities, and accentuate the integration between NLP and SE throughout the work. We make the key observation that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2023a; Luo et al., 2023), infilling objectives (Tay et al., 2023; Li et al., 2023g; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023h), architectural improvements (Shazeer, 2019; Su et al., 2021; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements are providing real-world testbeds for these technologies and driving the development of LLMs forward into production. We believe a systematic review of these advancements would benefit both communities.
@@
The rest of this work is organized following the taxonomy presented in Figure 1. In Section 2 we first provide the preliminaries of language modeling and Transformer models, and then in Section 3 we contextualize the evaluation of language models for code, highlighting the historical transition from various code understanding tasks to more practical text-to-code generation tasks. In Section 4 we discuss the plethora of LLMs that have demonstrated coding ability, and then in Section 5 we review the specialized and often smaller models by their architecture, with special attention on the recent application of infilling objectives, instruction tuning, reinforcement learning, and engineering improvements. Then, in Section 6, we discuss unique features of code that are not available to natural languages but have been utilized to aid code processing. In Section 7, we review the most recent integration between LLMs and software development, before finally concluding this work in Section 8 and highlighting the current challenges in code processing.
2 Background
In this section, we briefly review the preliminaries of Transformer-based language modeling, including common objectives for unidirectional and bidirectional models, as well as some popular models and designs in NLP.
2.1 Causal Language Modeling
Unidirectional language models (also known as causal language models222The training objective of such language models is Causal Language Modeling (CLM), but also referred to as Next Token Prediction.) factor the probability of a sentence into the product of each token’s conditional probability with the chain rule. A piece of input text consisting of tokens is modeled as
2.2 Masked Language Modeling
While this objective requires the model to have a deep understanding of the input text to reconstruct it, it suffers from low training efficiency, since only a small set of tokens (usually 15%) are masked (and thus “trained on”). To address this issue, Clark et al. (2020) propose ELECTRA, which is trained to discriminate whether or not each token in the input has been replaced by a BERT-like model instead.
2.3 Denoising Objectives
The vanilla Transformer encoder-decoder architecture combines the respective advantages of GPT and BERT. T5 (Raffel et al., 2020) is such a model pretrained with span corruption, which can be regarded as a variant of MLM. During pretraining, spans of text in the input are replaced with sentinel tokens, which plays the same role as [MASK] in BERT. The noisy input is first processed by the encoder with bidirectional attention, and the masked spans are then generated autoregressively by the decoder. Formally, if spans are sampled for corruption in input , the noisy input is then constructed by replacing each span with a special token <extra_id_i>, for , and the target is constructed by concatenating all spans prepended with corresponding sentinels: [<extra_id_1>, span, , <extra_id_k>, span]. The model is then trained with a standard seq2seq objective, by maximizing
Lester et al. (2021) show that models pretrained with such objectives can be adapted for autoregressive language modeling with extra pretraining using the prefix language modeling objective, i.e. spliting the text into two parts, processing the first part with encoder and generating the second part with decoder.
Tay et al. (2023) argue that span corruption is also closely related to CLM, since one can mask out the whole input text as a single span and train the decoder to generate it autoregressively. Inspired by this relation, they propose UL2, which is the combination of many span corruption objectives that differ in corruption rate and span length. Applying it to both encoder-decoder models and decoder-only models, they find that encoder-decoder models perform better under the same computation budget constraint. Other researches have also found that such encoder-decoder models generally perform better than causal decoder-only models (Wang et al., 2022b; Soltan et al., 2022).
2.4 Auxiliary Objectives
2.5 Implementation Design
PaLM (Chowdhery et al., 2022) introduces Rotary Position Embedding (RoPE) and Multi-Query Attention (MQA) into LLMs. RoPE (Su et al., 2021) multiplies the keys and queries of each self-attention layer by a position-dependent rotation matrix to inject position information, and is later shown to enable position interpolation for processing of longer sequences (Chen et al., 2023b; Rozière et al., 2023). Alternative to RoPE, Press et al. (2022) propose ALiBi, which directly attenuates the attention scores according to the relative position between key and query. This position embedding scheme is later adopted by BLOOM (Scao et al., 2022).
Apart from position embeddings, another issue in Transformer that has long troubled researchers is the fact that the complexity of self-attention scales quadratically with the input sequence length. Some works such as Reformer (Kitaev et al., 2020), Linformer (Wang et al., 2020c), Performer (Choromanski et al., 2021) and cosFormer (Qin et al., 2022b) use approximate attention to reduce this complexity, but they mostly come at the cost of degraded performance. Other works tackle this issue from an engineering point-of-view. MQA (Shazeer, 2019) shares the same set of keys and values across all attention heads to optimize memory-to-arithmetic ratio and significantly improves inference speed at small costs of model performance. Its variant Grouped-Query Attention (GQA, Ainslie et al., 2023) takes a middle-ground approach by dividing attention heads into groups and sharing the same set of keys/values within each group. Orthogonally, Dao et al. (2022) introduce FlashAttention, which is an exact but improved implementation of self-attention that optimizes IO operations on the accelerating device via tiling to improve memory efficiency.
Another important technique for improving LLMs’ training efficiency is mixed precision training (Micikevicius et al., 2018), which stores model weights and activations in lower precision to save memory consumption. Early works use FP16 format for this low precision, while BF16 has now become more popular (Chowdhery et al., 2022; Scao et al., 2022), as it can represent the same range of floating point numbers as FP32, thus preventing numerical overflowing and underflowing during training.
3 Evaluation of Language Models for Code
@@
Over the past decade, various evaluation tasks have been proposed by the software engineering community to evaluate code models. CodeXGLUE (Lu et al., 2021) consolidates most of such tasks into a single benchmark covering code understanding tasks such as clone detection, defect detection, and sequence-to-sequence generation tasks such as code repair, code translation, program synthesis, and code summarization. After Chen et al. (2021b) introduced HumanEval and Codex, however, text-to-code synthesis was brought into the spotlight in the NLP community and has since become a standard task for evaluating LLMs (Figure 3). Accordingly, we first briefly introduce each of the traditional tasks and the application of pretrained language models in them in Section 3.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in Section 3.2 and investigate program synthesis in more detail in Section 3.3. Lastly, we also discuss the latest trend of repository-level evaluation in Section 3.4. In Appendix A, we list benchmarks for each downstream task.
@@
[Code Evaluation
[Text-to-Code
@@
Retrieval
@@
RACS (Li et al., 2016), CodeMatcher (Liu et al., 2022a)
@@
TabCS (Xu et al., 2021), GraphSearchNet (Liu et al., 2023h),
TranCS (Sun et al., 2022)
TranS (Wang et al., 2020e), Heyman & Cutsem (2020),
@@
CCT-LM (Sorokin et al., 2023)
[Code
Synthesis
Euphony (Lee et al., 2018), Neo (Feng et al., 2018)
LPN (Ling et al., 2016), NSPS (Parisotto et al., 2017),
DeepCoder (Balog et al., 2017), RobustFill (Devlin et al., 2017a),
(Yin & Neubig, 2017), ASN (Rabinovich et al., 2017),
NGDS (Kalyan et al., 2018), Bunel et al. (2018),
ReCode (Hayati et al., 2018), AutoPandas (Bavishi et al., 2019),
Wei et al. (2019), PlotCoder (Chen et al., 2021c),
Huang et al. (2022a)
TreeGen (Sun et al., 2020), REDCODER (Parvez et al., 2021),
Jigsaw (Jain et al., 2022), JuPyT5 (Chandel et al., 2022),
CodeT (Chen et al., 2023a), TiCoder (Lahiri et al., 2022),
AceCoder (Li et al., 2023d), Self-Debugging (Chen et al., 2023c),
ClarifyGPT (Mu et al., 2023)
[Text-to-SQL
@@
GNN (Bogin et al., 2019), TREQS (Wang et al., 2020b)
@@
Graphix-T5 (Li et al., 2023e),
RESDSQL (Li et al., 2023b), Liu et al. (2023a),
Self-Debugging (Chen et al., 2023c), DIN-SQL (Pourreza & Rafiei, 2023),
@@
Guo et al. (2023a), DAIL-SQL (Gao et al., 2023a)
[Math Programming
@@
GSM8K-Python (Chowdhery et al., 2022; Wang et al., 2023d)
[Code-to-Code]
[Code-to-Text]
[Code-to-Pattern]
[Text-to-Text]
3.1 Downstream Tasks of Code Processing
Following the custom in software engineering, we categorize the evaluation tasks for code according to their input/output modality, and break down these tasks into five families: text-to-code, code-to-code, code-to-text, code-to-pattern, and text-to-text. We note that this taxonomy is interleaved with the understanding-generation dichotomy in NLP, since each category may contain both understanding and generation tasks, as discussed in Section 3.1.6.

### v5

@@ Abstract
In this work we systematically review the recent advancements in code processing with language models, covering 50+ models, 30+ evaluation tasks, 170+ datasets, and 800 related works. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also discuss code-specific features such as AST, CFG, and unit tests, along with their application in training code language models, and identify key challenges and potential future directions in this domain.
@@ 1 Introduction
Language modeling has advanced remarkably in recent years with the advent of pretrained Transformers (Vaswani et al., 2017) such as BERT (Devlin et al., 2019) and GPT (Radford et al., 2018). As large language models (LLMs) scaled to hundreds of billions of parameters and started to display early signs of artificial general intelligence (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023), their applications have also transcended text processing. Pioneered by Codex (Chen et al., 2021b), LLMs have achieved impressive results in code processing, giving rise to commercial products such as GitHub Copilot111https://github.com/features/copilot and open-source multi-billion code models such as StarCoder (Li et al., 2023h) and Code LLaMA (Rozière et al., 2023).
The application of pretrained Transformers in code processing, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for code in this work, covering 50+ models, 30+ downstream tasks, 170+ datasets, and 800 related works. We break down different categories of code language models, ranging from colossal models trained on general domains to tiny models trained specifically for code understanding or generation. We emphasize on the relations and differences between such models, and highlight the integration of code-specific features, such as abstract syntax trees or data flows, into language models, as well as the latest techniques adapted from NLP.
Related to our work, we are aware of several surveys on similar topics, with three works concurrent to us (Hou et al., 2023; Zheng et al., 2023b; She et al., 2023). These works, however, focus either on NLP side (Zan et al., 2023; Xu & Zhu, 2022) or SE side (Niu et al., 2023; Hou et al., 2023; Zheng et al., 2023b; She et al., 2023), and do not cover models, tasks, and challenges from the other side. For example, Zan et al. (2023) focus on LLMs for text-to-code generation, while giving little discussion of other applications in software engineering community. Hou et al. (2023) and She et al. (2023), in contrast, comprehensively review works from SE venues such as ASE and ICSE, but cite only a handful of works from deep learning and NLP venues such as ACL, EMNLP, NeurIPS, and ICLR.
@@
where level=2text width=5.2em,font=,
@@
[Code LMs, fill=black, fill opacity=0.7, text=white
[Raw LM, fill=Goldenrod, fill opacity=0.4
[LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), GPT-NeoX (Black et al., 2022),
@@
LLaMA 2 (Touvron et al., 2023b), Phi-1.5 (Li et al., 2023i), Baichuan 2 (Yang et al., 2023a),
Qwen (Bai et al., 2023), Mistral (Jiang et al., 2023a), Gemini (Anil et al., 2023a),
DeepSeek (DeepSeek-AI et al., 2024),
Mixtral (Jiang et al., 2024a),
Gemma (Mesnard et al., 2024),
Yi (Young et al., 2024) …,
text width=27.7em,font=, fill=Goldenrod, fill opacity=0.4]
[Adapted LM, fill=Apricot, fill opacity=0.4
[Codex (Chen et al., 2021b), PaLM Coder (Chowdhery et al., 2023), Minerva (Lewkowycz et al., 2022),
PaLM 2* (Anil et al., 2023b), Code LLaMA (Rozière et al., 2023), Code-Qwen (Bai et al., 2023),
text width=27.7em,font=, fill=Apricot, fill opacity=0.4]
[Specialized LM, fill=MidnightBlue, fill opacity=0.4
[Encoder, fill=ProcessBlue, fill opacity=0.4
@@
CodeSage (Zhang et al., 2024),
text width=20.5em, fill=ProcessBlue, fill opacity=0.4]
[Decoder, fill=Cerulean, fill opacity=0.4
[CLM, fill=Cerulean, fill opacity=0.4
@@
text width=16.8em, fill=Cerulean, fill opacity=0.4]
[FIM, fill=Cerulean, fill opacity=0.4
[InCoder (Fried et al., 2023), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023h),
DeepSeek Coder (Guo et al., 2024),
StarCoder 2 (Lozhkov et al., 2024),
CodeShell (Xie et al., 2024),
text width=16.8em, fill=Cerulean, fill opacity=0.4]
[UniLM, fill=Cyan, fill opacity=0.4
@@
text width=20.5em, fill=Cyan, fill opacity=0.4]
[Encoder-Decoder, fill=CornflowerBlue, fill opacity=0.4
@@
CodeT5+ (Wang et al., 2023f),
AST-T5 (Gong et al., 2024),
text width=20.5em, fill=CornflowerBlue, fill opacity=0.4]
@@
Finetuning, fill=WildStrawberry, fill opacity=0.4
[Instruction Finetuning, fill=RedOrange, fill opacity=0.4
[WizardCoder (Luo et al., 2023), PanGu-Coder2 (Shen et al., 2023), OctoCoder (Muennighoff et al., 2024),
MFTCoder (Liu et al., 2023b), WaveCoder (Yu et al., 2023c),
Astraios (Zhuo et al., 2024),
CCT (Jiang et al., 2024b),
text width=20.5em, fill=RedOrange, fill opacity=0.4]
[Reinforcement Learning, fill=Bittersweet, fill opacity=0.4
@@
StepCoder (Dou et al., 2024),
text width=20.5em, fill=Bittersweet, fill opacity=0.4]
@@
Thus, building on these works, we endeavor to unite the perspectives from both communities, and accentuate the integration between NLP and SE throughout the work. We make the key observation that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2024; Luo et al., 2023), infilling objectives (Tay et al., 2023b; Li et al., 2023h; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023i), architectural improvements (Shazeer, 2019; Su et al., 2024; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements are providing real-world testbeds for these technologies and driving the development of LLMs forward into production. We believe a systematic review of these advancements would benefit both communities.
@@
The rest of this work is organized following the taxonomy presented in Figure 1. In Section 2 we first contextualize the downstream tasks in code processing, highlighting the historical transition from various code understanding tasks to more practical text-to-code generation tasks. Then, in Section 3 we provide the preliminaries of language modeling and Transformer models, and in Section 4 we discuss the plethora of LLMs that have demonstrated coding ability. In Section 5 we review the specialized and often smaller models by their architecture, with special attention on the recent application of infilling objectives, instruction tuning, reinforcement learning, and engineering improvements. Then, in Section 6, we discuss unique features of code that are not available to natural languages but have been utilized to aid code processing. In Section 7, we review the most recent integration between LLMs and software development, before finally concluding this work in Section 8 and highlighting the current challenges in code processing.
2 Downstream Tasks in Code Processing
@@
Over the past decade, the software engineering community has found applications of code models in various downstream tasks. CodeXGLUE (Lu et al., 2021) consolidates many of such tasks into a single benchmark covering code understanding tasks such as clone detection, defect detection, and sequence-to-sequence generation tasks such as code repair, code translation, program synthesis, and code summarization. After Chen et al. (2021b) introduced HumanEval and Codex, however, text-to-code synthesis was brought into the spotlight in the NLP community and has since become a standard task for evaluating LLMs (Figure 2). Accordingly, we first briefly introduce each of the traditional tasks and the application of pretrained language models in them in Section 2.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in Section 2.2 and investigate program synthesis in more detail in Section 2.3. Lastly, we also discuss the latest trend of repository-level coding in Section 2.4. In Appendix A, we list benchmarks for each downstream task.
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5
@@
Retrieval, fill=yellow, fill opacity=0.5
@@
RACS (Li et al., 2016), CodeMatcher (Liu et al., 2022a), fill=yellow, fill opacity=0.5
@@
TabCS (Xu et al., 2021a), GraphSearchNet (Liu et al., 2023i),
TranCS (Sun et al., 2022), fill=yellow, fill opacity=0.5
TranS3 (Wang et al., 2020e), Heyman & Cutsem (2020),
@@
CCT-LM (Sorokin et al., 2023), fill=yellow, fill opacity=0.5
[Text-to-SQL, fill=yellow, fill opacity=0.5
[, fill=yellow, fill opacity=0.5
@@
GNN (Bogin et al., 2019), TREQS (Wang et al., 2020b), fill=yellow, fill opacity=0.5
@@
Graphix-T5 (Li et al., 2023f),
RESDSQL (Li et al., 2023c), Liu et al. (2023a),
Self-Debugging (Chen et al., 2023d), DIN-SQL (Pourreza & Rafiei, 2023),
@@
Guo et al. (2023a), DAIL-SQL (Gao et al., 2023a),
SQL-Encoder (Pourreza et al., 2024), fill=yellow, fill opacity=0.5
[Code
Synthesis, fill=yellow, fill opacity=0.5
Euphony (Lee et al., 2018), Neo (Feng et al., 2018), fill=yellow, fill opacity=0.5
LPN (Ling et al., 2016), NSPS (Parisotto et al., 2017),
DeepCoder (Balog et al., 2017), RobustFill (Devlin et al., 2017a),
(Yin & Neubig, 2017), ASN (Rabinovich et al., 2017),
NGDS (Kalyan et al., 2018), Bunel et al. (2018),
ReCode (Hayati et al., 2018), AutoPandas (Bavishi et al., 2019),
Wei et al. (2019), PlotCoder (Chen et al., 2021c),
Huang et al. (2022a), fill=yellow, fill opacity=0.5
TreeGen (Sun et al., 2020), REDCODER (Parvez et al., 2021),
Jigsaw (Jain et al., 2022), JuPyT5 (Chandel et al., 2022),
CodeT (Chen et al., 2023a), TiCoder (Lahiri et al., 2022),
AceCoder (Li et al., 2023e), Self-Debugging (Chen et al., 2023d),
ClarifyGPT (Mu et al., 2023),
MANGO (Chen et al., 2024), fill=yellow, fill opacity=0.5
[Math Programming, fill=yellow, fill opacity=0.5
[, fill=yellow, fill opacity=0.5
[, fill=yellow, fill opacity=0.5
@@
GSM8K-Python (Chowdhery et al., 2023; Wang et al., 2023f), fill=yellow, fill opacity=0.5
[Code-to-Code, fill=cyan, fill opacity=0.4]
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2]
[Text-to-Text, fill=Dandelion, fill opacity=0.4]
2.1 Downstream Applications of Code Modeling
Following the custom in software engineering, we categorize the applications of code models according to their input/output modality, and break down these tasks into five families: text-to-code, code-to-code, code-to-text, code-to-pattern, and text-to-text. We note that this taxonomy is interleaved with the understanding-generation dichotomy in NLP, since each category may contain both understanding and generation tasks, as discussed in Section 2.1.6.